Hi i need to find the solution on this problem:
the product of two number is 30 and their difference is 7. find the numbers.

Answer by stanbon(75887)

the product of two number is 30 and their difference is 7. find the numbers.
---------
Let one of the numbers be "x"; the other number is "x-7"
EQUATION:
x(x-7)=30
x^2-7x-30=0
(x-10)(x+3)=0
x=10 and x-7=3
OR
x=-3 and x-7=-10
===================================
Cheers,
Stan H.